The price elasticity of demand for chicken is estimated to be -0.45. If the price of chicken Increased by 08 percent, what will be the expected percentage decrease in the quantity of chicken sold.


1
Expert's answer
2020-06-23T04:26:58-0400

The price elasticity of demand is given by:



"E_d = \dfrac{\text{\% Change in quantity demanded}}{\text{\% Change in price}}"

If the elasticity is -0.45 and the price increases by 8%, then:



"-0.45 = \dfrac{\text{\% Change in quantity demanded}}{\text{8\%}}\\[0.3cm]
\text{\% Change in quantity demanded} = -0.45\times 8\%\\[0.3cm]
\color{red}{\text{\% Change in quantity demanded} = -3.6\%}"

The quantity demanded will decrease by 3.6%.